OpenStudy (anonymous):

What is the equation of the line of symmetry for the graph of y = 3(x + 2)2 - 4?

OpenStudy (anonymous):

Write the equation in vertex form. y = 3x² + 12x - 2 = 3(x² + 4x) - 2 = 3(x² + 4x + (4/2)² - (4/2)²) - 2 = 3(x² + 4x + 4) - 3*4 - 2 = 3(x + 2)² - 14

OpenStudy (anonymous):

Answer= -14

OpenStudy (anonymous):

a. x = -4 b. x = -2 c. x = 2 d. x = 3

ganeshie8 (ganeshie8):

Hint : line of symmetry goes thru vertex

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

y = 3x² + 12x - 2 = 3(x² + 4x) - 2 = 3(x² + 4x + (4/2)² - (4/2)²) - 2 = 3(x² + 4x + 4) - 3*4 - 2 = 3(x + 2)² - 14 The equation for the axis of symmetry is (x + 2) = 0, that is, x = -2. The axis passes through the vertex of the parabola; its coordinates are (-2, -14)